OpenStudy (anonymous):

a trip to portland,oragan, from boston will take 7 3/4 hours. assuming we are two thirds of the way there, how much longer, in hours, will the trip take?

OpenStudy (perl):

it might help to change 7 3/4 to an improper fraction

OpenStudy (perl):

7 3/4 = 7 + 3/4 = 7/1 + 3/4 = 28/4 + 3/4 = 31/4

OpenStudy (perl):

now the question says, assuming we are 2/3 of the way , how much time is left for the trip. If you are 2/3 of the way, that leaves you with 1/3 of the way left. So multiply 1/3 and 31/4